Consider a monopoly whose total cost function is TC=Q3-30Q2+302Q, whose marginal cost function is MC= 3Q2-60Q+302, whose demand is P=329-30Q, and whose marginal revenue function MR= 329-60Q, where Q is output and P is price.

Assume that the firm maximizes profit but cannot practice price discrimination 


How much does the firm produce?        

How much does the firm charge?

How large are the firms profit?


1
Expert's answer
2021-06-20T17:51:28-0400

a) Since MR= MC, then "3Q^2-60Q + 302 = 329 -60Q"

"3Q^2 = 27"

"Q^2 = 27""\/3"

"Q^2 = 9"

"Q = 3"

b)

"P =329-30Q"

"P = 329 - 30(3)"

"P = 329 -90"

"P = 239"

C) Considering "Profit = PQ - TC""Q = 3 and P = 239"

"TC=Q^3-30Q^2+302Q"

"= 3^3 -30 (3^2) + 302(3)"

"= 27 -270 + 906"

"TC = 663"

"PQ = 239 \u00d7 3 = 717"

"= 717 -663"

"= 54"